Return on equity is not 27.35% or 27.36%. Please write the steps to find what the correct return on equity should be. Every other step is already correct. The 2021 income statement for Anderson TV and Appliance reported net sales of $230,000 and net income of $90,000. Average tot
Answer #1

Ending stockholder's equity = 300000+90000-30000 = 360000

Average equity = (300000+360000) /2= 330000

return on equity = return / average equity = 90000/330000*100= 27.27%